# Reconstruction-based Multi-Normal Prototypes Learning for Weakly Supervised Anomaly Detection

### v1

@@ Abstract
Anomaly detection is a crucial task in various domains. Most of the existing methods assume the normal sample data clusters around a single central prototype while the real data may consist of multiple categories or subgroups. In addition, existing methods always assume all unlabeled data are normal while they inevitably contain some anomalous samples. To address these issues, we propose a reconstruction-based multi-normal prototypes learning framework that leverages limited labeled anomalies in conjunction with abundant unlabeled data for anomaly detection. Specifically, we assume the normal sample data may satisfy multi-modal distribution, and utilize deep embedding clustering and contrastive learning to learn multiple normal prototypes to represent it. Additionally, we estimate the likelihood of each unlabeled sample being normal based on the multi-normal prototypes, guiding the training process to mitigate the impact of contaminated anomalies in the unlabeled data. Extensive experiments on various datasets demonstrate the superior performance of our method compared to state-of-the-art techniques.
1 Introduction
@@
Anomaly detection aims to identify samples that deviate significantly from the general data distribution, known as anomalies or outliers (Jiang et al. 2023). These anomalies often indicate potential problems or noteworthy events that require attention (Pang et al. 2021). The significance of anomaly detection spans a variety of critical domains, such as network security (Estévez-Pereira, Fernández, and Novoa 2020), Internet of Thing (IoT) security (Bulla and Birje 2022), and financial fraud detection (Popat and Chaudhary 2018). In each of these areas, effective anomaly detection methods are essential due to the substantial impact of anomalies.
@@
In real-world scenarios, obtaining labeled data for anomaly detection is often challenging and costly (Jiang et al. 2023, 2024). Consequently, fully-supervised anomaly detection is impractical. While unsupervised methods bypass the need for extensive labeling, they struggle to achieve optimal performance without knowing what true anomalies look like. This limitation frequently results in the misidentification of noisy or irrelevant data points as anomalies (Pang et al. 2023). To address these challenges, weakly supervised anomaly detection methods have emerged (Pang, Shen, and van den Hengel 2019; Ruff et al. 2020; Zhou et al. 2021; Xu et al. 2023b). These methods leverage a small portion of labeled anomalous data in conjunction with a large amount of unlabeled data, primarily composed of normal instances. This kind of approach is particularly significant in practical applications where we can only label a small fraction of anomalies.
Figure 1: Motivations of multi-normal prototypes learning.
@@
However, existing anomaly detection methods still face two major challenges. First, most of the existing anomaly detection methods (Ruff et al. 2018a, 2020; Pang, Shen, and van den Hengel 2019) assume that normal data clusters around a single central prototype. However, real-world data often consists of multiple categories or subgroups (Fu, Zhang, and Fan 2024), making this assumption overly simplistic, as shown in Figure 1. Second, existing methods always assume all unlabeled data are normal while they inevitably contain some anomalous samples. Due to the assumption, many studies (Zhou et al. 2021; Fu, Zhang, and Fan 2024; Pang, Shen, and van den Hengel 2019) directly employ large amounts of unlabeled data as inliers for model training to identify the distribution of normal data patterns. These methods are vulnerable to the presence of occasional anomalies (i.e., anomaly contamination) within the unlabeled data, and the detection accuracy rapidly declines as the proportion of mixed-in anomalies increases (Pang et al. 2023; Xu et al. 2023b).
@@
To address the problems stated above, we propose a reconstruction-based multi-normal prototype learning framework for weakly supervised anomaly detection. Different from existing reconstruction-based anomaly detection methods (Zhou et al. 2021), we treat the normal samples and anomalous samples differently during the latent representation learning with consideration of the likelihood of each unlabeled sample being normal. To better estimate the likelihood of normal and detect anomalies, we propose to learn multiple normal prototypes in the latent space with deep embedding clustering and contrastive learning. Finally, we compute a comprehensive anomaly score with consideration of both the information of sample reconstruction and multiple normal prototypes.
@@
  • We propose a novel anomaly detection framework that combines reconstruction learning with multi-normal prototype learning. Extensive experiments across various datasets demonstrate that our method outperforms state-of-the-art techniques.
@@
  • We propose to estimate and take into consideration the likelihood of each unlabeled sample being normal during model training, which can enhance resistance to anomaly contamination and more effectively detect unseen anomaly classes combined with previous designs.
2 Related Work
2.1 Anomaly Detection Based on Prototype Learning
One-class classification methods (Ruff et al. 2018b) are a fundamental category of traditional anomaly detection techniques. These methods are trained solely on normal data and identify anomalies by detecting deviations from a central prototype. Examples include One-Class SVM (OC-SVM) (Schölkopf et al. 1999), Support Vector Data Description (SVDD) (Tax and Duin 2004), Deep SVDD (Ruff et al. 2018a), and DeepSAD (Ruff et al. 2020), all of which assume that normal samples cluster around a single prototype in feature space, with anomalies located at the periphery. However, some studies have shown that normal data frequently consists of multiple categories or subgroups (Fu, Zhang, and Fan 2024), rendering a single prototype insufficient. To address these limitations, multi-prototype learning approaches have been developed to better represent the diversity within normal data, thereby enhancing anomaly detection capabilities. This concept has been applied across various domains, such as video anomaly detection (Park, Noh, and Ham 2020), image anomaly detection (Huang, Kang, and Wu 2024), and time series analysis (Li, Jentsch, and Müller 2023). These advancements highlight the potential of multi-prototype learning, particularly in the context of tabular data for weakly supervised anomaly detection.
@@ 2.1 Anomaly Detection Based on Prototype Learning
2.2 Weakly Supervised Anomaly Detection
@@
Weakly supervised anomaly detection has emerged as a solution to the limitations of purely unsupervised methods, which struggle to identify what anomalies look like due to the lack of labeled examples (Pang et al. 2023). In weakly supervised settings, a small number of labeled anomalies are combined with a large amount of unlabeled data to enhance detection performance. Existing weakly supervised methods face notable challenges. Many rely heavily on large amounts of unlabeled data to identify normal data patterns, making them susceptible to contamination by occasional anomalies within the unlabeled dataset (Zhou et al. 2021; Ruff et al. 2020). This lack of effective mechanisms to mitigate the impact of these anomalies can significantly degrade detection performance, with accuracy declining rapidly as the proportion of mixed-in anomalies increases (Zhou et al. 2021; Fu, Zhang, and Fan 2024).
3 Methodology
3.1 Task Definition
@@
In the context of weakly supervised anomaly detection, we are given a dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N+K}$, which contains two disjoint sets, $\mathcal{X_{U}}$ and $\mathcal{X_{A}}$. Here, $\mathcal{X_{U}}=\{x_{i}\}_{i=1}^{N}$ represents a large set of unlabeled samples, and $\mathcal{X_{A}}=\{x_{i}\}_{i=N+1}^{N+K}$ (where $K\ll N$) denotes a small set of labeled anomaly samples. Our objective is to learn a scoring function $\phi:\mathcal{X}\rightarrow\mathbb{R}$ that assigns anomaly scores to each data instance, with the larger value of $\phi(x_{i})$ indicating a higher likelihood of $x_{i}$ being an anomaly.
3.2 Overview of Our Method
@@
To detect anomalies in a weakly supervised setting, we propose a novel approach combining reconstruction learning and multi-normal prototype learning, as illustrated in Figure 2. Our method consists of three main components: 1) The Reconstruction Learning Module aims to guide the latent representation learning with differentiated treatment to reconstruction errors of normal and anomalous samples; 2) The Multi-Normal Prototypes Learning Module aims to model the distribution of normal data better and utilizes deep embedding clustering and contrastive learning to build multiple normal prototypes for anomaly detection; 3) The Unified Anomaly Scoring Module aims to compute a comprehensive anomaly score with consideration of both the information of sample reconstruction and multiple normal prototypes. The detailed implementation of each component is described below.
3.3 Reconstruction Learning Module
This module serves two primary purposes: 1) transforming the data into a latent space to capture the essential features of samples; and 2) differentiating between normal and anomalous samples through reconstruction errors.
@@ 3.3 Reconstruction Learning Module
The decoder function $\psi^{\prime}$ maps the hidden representation $z_{i}$ back to a reconstruction $\hat{x}_{i}\in\mathbb{R}^{D}$ of the original input:
@@
On the other hand, to distinguish normal samples from anomalies, we apply a hinge loss (Cortes and Vapnik 1995) to ensure that normal samples have a lower reconstruction loss compared to anomalies, maintaining a sufficient margin $m_{1}$. Therefore, the reconstruction loss for labeled anomalies is defined as:
$\mathcal{L_{\text{recon}}^{\text{a}}}=\mathbb{E}_{x_{i}\sim\mathcal{X_{A}}}[\max\left(0,m_{1}-(\mathit{l}(x_{i},\hat{x}_{i})-L_{recon}^{u})\right)]$ (4)
@@
3.4 Multi-Normal Prototypes Learning Module
@@
Initialization and Similarity Calculation.
At the beginning of training, we initialize $k$ normal prototypes by clustering normal data’s latent representations from the pre-train Encoder following (Guo et al. 2017), denoted as $\{u_{j}\in\mathbb{R}^{H}\}_{j=1}^{k}$. During the model training and inference phases, we need to calculate the similarity between a given sample $x_{i}$ and each prototype $u_{j}$ to determine which category or subgroup the sample belongs to. Specifically, we assume the samples of each category or subgroup satisfy the Student’s $t$-distribution, and calculate the similarity between sample $x_{i}$ and prototype $u_{j}$ as follows:
$s_{ij}=Sim(z_{i};\Theta_{\{u_{j}\}_{j=1}^{k}})=\left(1+\frac{\|z_{i}-u_{j}\|^{2}}{\alpha}\right)^{-\frac{\alpha+1}{2}}$ (6)
where $z_{i}$ denotes the latent representation of sample $x_{i}$, $\alpha$ is the degree of freedom of the Student’s t-distribution, $\Theta_{\{u_{j}\}_{j=1}^{k}}$ is the parameters of the multi-normal prototypes, consistent with the neural network parameters that can be optimized via backpropagation. To ensure consistency and stability during training, both the normal prototype $u_{j}$ and the latent representation $z_{i}$ are normalized before calculating the similarity.
Weights of Unlabeled Samples.
@@
Clustering and Contrastive Learning.
@@
First, recognizing that the relationship between normal samples and multiple normal prototypes aligns with an unsupervised clustering problem, we employ Deep Embedding Clustering (Xie, Girshick, and Farhadi 2016), which utilizes deep neural networks to simultaneously learn feature representations and cluster assignments. This process involves optimizing a KL divergence loss for unlabeled samples:
@@
Table 1: Overview of Datasets. $D$ and $N$ denote data dimension and size, respectively. $\delta$ indicates the anomaly ratio.

### v2

# Multi-Normal Prototypes Learning for Weakly Supervised Anomaly Detection

@@ Abstract
Anomaly detection is a crucial task in various domains. Most of the existing methods assume the normal sample data clusters around a single central prototype while the real data may consist of multiple categories or subgroups. In addition, existing methods always assume all unlabeled samples are normal while some of them are inevitably being anomalies. To address these issues, we propose a novel anomaly detection framework that can efficiently work with limited labeled anomalies. Specifically, we assume the normal sample data may consist of multiple subgroups, and propose to learn multi-normal prototypes to represent them with deep embedding clustering and contrastive learning. Additionally, we propose a method to estimate the likelihood of each unlabeled sample being normal during model training, which can help to learn more efficient data encoder and normal prototypes for anomaly detection. Extensive experiments on various datasets demonstrate the superior performance of our method compared to state-of-the-art methods. Our codes are available at: https://github.com/Dongzhijin/MNPWAD
Index Terms:
Anomaly Detection, Multi-normal Prototypes, Anomaly Contamination, Weakly Supervised Learning.
I Introduction
@@
Anomaly detection aims to identify samples that deviate significantly from the general data distribution, known as anomalies or outliers [1]. These anomalies often indicate potential problems or noteworthy events that require attention [2]. The significance of anomaly detection spans a variety of critical domains, such as network security [3], Internet of Thing (IoT) security [4], and financial fraud detection [5]. In each of these areas, effective anomaly detection methods are essential due to the substantial impact of anomalies, which can lead to severe consequences such as security breaches[6], financial losses[7], compromised privacy[8], and even risks to human health and safety[9, 10].
@@
In real-world scenarios, obtaining labeled data for anomaly detection is often challenging and costly [1, 11, 12]. Although unsupervised methods[13, 14] can bypass the need for extensive labeling, they struggle to achieve optimal performance without knowing what true anomalies look like. This limitation frequently results in the misidentification of many normal data points as anomalies [15]. To address these challenges, weakly supervised anomaly detection methods have emerged [16, 17]. These methods leverage a small portion of labeled anomalous data in conjunction with a large amount of unlabeled data, primarily composed of normal instances. This kind of approach is particularly significant in practical applications where we can only label a small fraction of anomalies.
Figure 1: Motivations of the proposed multi-normal prototype learning (b). Compared with single-normal prototype learning (a), multi-normal prototype (b) can more effectively capture the diversity of normal data and estimate the probability of unlabeled samples being normal to resist anomaly contamination.
@@
However, existing anomaly detection methods still face several significant challenges. First, most of the existing anomaly detection methods [18, 19, 16] assume that normal data clusters around a single central prototype. This assumption oversimplifies real-world scenarios where normal data often comprises multiple categories or subgroups, as illustrated in Fig. 1. Such methods struggle to capture the complexity and diversity of normal data, leading to potential misclassification of normal samples as anomalies [20]. Second, existing methods always assume all unlabeled data are normal while they inevitably contain some anomalous samples. Due to the assumption, many studies [21, 20, 16] directly employ large amounts of unlabeled data as inliers for model training to identify the distribution of normal data patterns. These methods are vulnerable to the presence of occasional anomalies (i.e., anomaly contamination)[22] within the unlabeled data, and the detection accuracy rapidly declines as the proportion of mixed-in anomalies increases [15, 17].
@@
To address the problems stated above, we propose a reconstruction-based multi-normal prototype learning framework for weakly supervised anomaly detection. Different from existing reconstruction-based anomaly detection methods [21], we treat the normal samples and anomalous samples differently during the latent representation learning with consideration of the likelihood of each unlabeled sample being normal. To better estimate the likelihood of normal and detect anomalies, we propose to learn multiple normal prototypes in the latent space with deep embedding clustering and contrastive learning. Finally, we compute a comprehensive anomaly score with consideration of both the information of sample reconstruction and multiple normal prototypes.
@@
  • We propose a novel anomaly detection framework that combines reconstruction learning with multi-normal prototype learning. Extensive experiments across various datasets demonstrate that our method significantly outperforms state-of-the-art methods.
@@
  • We propose to estimate and take into consideration the likelihood of each unlabeled sample being normal during model training, which can enhance resistance to anomaly contamination and more effectively detect unseen anomaly classes.
II Related Work
II-A Anomaly Detection Based on Prototype Learning
One-class classification methods [23] are a fundamental category of traditional anomaly detection techniques. These methods are trained solely on normal data and identify anomalies by detecting deviations from a central prototype. Examples include One-Class SVM (OC-SVM) [24], Support Vector Data Description (SVDD) [25], Deep SVDD [18], and DeepSAD [19], all of which assume that normal samples cluster around a single prototype in feature space, with anomalies located at the periphery. However, some studies have shown that normal data frequently consists of multiple categories or subgroups [20], rendering a single prototype insufficient [26]. To address these limitations, multi-prototype learning approaches have been developed to better represent the diversity within normal data, thereby enhancing anomaly detection capabilities. This concept has been applied across various domains, such as video anomaly detection [27], image anomaly detection [28], and time series analysis [29, 30, 31]. These advancements highlight the potential of multi-prototype learning, particularly in the context of tabular data for weakly supervised anomaly detection. By capturing the diversity inherent in normal data through multiple prototypes, our proposed method aims to more effectively detect anomalies that would otherwise be missed by single-prototype models.
@@ II-A Anomaly Detection Based on Prototype Learning
II-B Weakly Supervised Anomaly Detection
@@
Weakly supervised anomaly detection has emerged as a solution to the limitations of purely unsupervised methods, which struggle to identify what anomalies look like due to the lack of labeled examples [15]. In weakly supervised settings, a small number of labeled anomalies are combined with a large amount of unlabeled data to enhance detection performance. Existing weakly supervised methods face notable challenges. Many rely heavily on large amounts of unlabeled data to identify normal data patterns, making them susceptible to contamination by occasional anomalies within the unlabeled dataset [21, 19]. This lack of effective mechanisms to mitigate the impact of these anomalies can significantly degrade detection performance, with accuracy declining rapidly as the proportion of mixed-in anomalies increases [21, 20]. Moreover, a common issue across these methods [15, 17, 19, 16] is overfitting to the limited labeled anomalies. This over-fitting weakens the generalization capability of the models, especially when the number of known anomalies is small or when the models are confronted with previously unseen anomaly types during the testing phase. As a result, detection performance suffers, highlighting the need for methods that can better leverage the small amount of labeled data while effectively utilizing the distribution patterns within a large pool of unlabeled data.
III Methodology
III-A Task Definition
@@
Let $\mathcal{X}=\{x_{i}\}_{i=1}^{N+K}$ denote the sample set, which contains two disjoint subsets, $\mathcal{X_{U}}$ and $\mathcal{X_{A}}$. Here, $\mathcal{X_{U}}=\{x_{i}\}_{i=1}^{N}$ represents a large set of unlabeled samples, and $\mathcal{X_{A}}=\{x_{i}\}_{i=N+1}^{N+K}$ (where $K\ll N$) denotes a small set of labeled anomaly samples. Our objective is to learn a scoring function $\phi:\mathcal{X}\rightarrow\mathbb{R}$ that assigns anomaly scores to each data instance, with the larger value of $\phi(x_{i})$ indicating a higher likelihood of $x_{i}$ being an anomaly.
III-B Overview of Our Method
@@
To detect anomalies in a weakly supervised setting, we propose a novel approach combining reconstruction learning and multi-normal prototype learning, as illustrated in Fig. 2. It consists of three main components: 1) Reconstruction Learning Module aims to guide the latent representation learning with differentiated treatment to reconstruction errors of normal and anomalous samples; 2) Multi-Normal Prototypes Learning Module aims to model the distribution of normal data better and utilizes deep embedding clustering and contrastive learning to build multiple normal prototypes for anomaly detection; 3) Unified Anomaly Scoring Module aims to compute a comprehensive anomaly score with consideration of both the information of sample reconstruction and multiple normal prototypes. The detailed design of each component is described below.
III-C Reconstruction Learning Module
The primary purpose of this module is to transform the data into a latent space that captures the essential features of each sample, enabling effective differentiation between normal and anomalous samples. These latent representations are subsequently utilized in the Multi-Normal Prototypes Learning Module and serve as one of the inputs to the Unified Anomaly Scoring Module. Unlike traditional reconstruction-based methods[32, 33], which aim to minimize the reconstruction error for all samples without considering their labels, our approach incorporates label information and the likelihood of unlabeled samples being normal.
@@ III-C Reconstruction Learning Module
The decoder function $\psi^{\prime}$ maps the latent representation $z_{i}$ back to a reconstruction $\hat{x}_{i}\in\mathbb{R}^{D}$ of the original input:
@@
On the other hand, to distinguish normal samples from anomalies, we apply a hinge loss [34] to ensure that normal samples have a lower reconstruction loss compared to anomalies, maintaining a sufficient margin $m_{1}$. Therefore, the reconstruction loss for labeled anomalies is defined as:
$\mathcal{L_{\text{recon}}^{\text{a}}}=\mathbb{E}_{x_{i}\sim\mathcal{X_{A}}}[\max\left(0,m_{1}-(\mathit{l}(x_{i},\hat{x}_{i})-\mathcal{L_{\text{recon}}^{\text{u}}})\right)]$ (4)
@@
III-D Multi-Normal Prototypes Learning Module
@@
Initialization and Similarity Calculation
At the beginning of training, we initialize $k$ normal prototypes by clustering normal data’s latent representations from the pre-train Encoder, denoted as $\{u_{j}\in\mathbb{R}^{H}\}_{j=1}^{k}$. During the model training and inference phases, we need to calculate the similarity between a given sample $x_{i}$ and each prototype $u_{j}$ to determine which category or subgroup the sample belongs to. Specifically, we assume the samples of each category or subgroup satisfy the Student’s $t$-distribution, and calculate the similarity between sample $x_{i}$ and prototype $u_{j}$ as follows:
$s_{ij}=\left(1+\frac{\|z_{i}-u_{j}\|^{2}}{\alpha}\right)^{-\frac{\alpha+1}{2}}$ (6)
where $z_{i}$ denotes the latent representation of sample $x_{i}$, $\alpha$ is the degree of freedom of the Student’s t-distribution, $u_{j}$ denotes the latent representation of the $j$-th normal prototype. To ensure consistency and stability during training, both the normal prototype $u_{j}$ and the latent representation $z_{i}$ are normalized before calculating the similarity.
Weights of Unlabeled Samples
@@
Clustering and Contrastive Learning
@@
First, recognizing that the relationship between normal samples and multiple normal prototypes aligns with an unsupervised clustering problem, we employ Deep Embedding Clustering [35] to simultaneously learn feature representations and cluster assignments. This process involves optimizing a KL divergence loss for unlabeled samples:
@@
III-E Unified Anomaly Scoring Module
To detect anomalies with consideration of both the reconstruction error $e_{i}=\mathit{l}(x_{i},\hat{x}_{i})$ and the multi-normal prototype information, we design a unified anomaly scoring module. Inspired by the residual network [36], we also take the latent representation $z_{i}$ of the given sample $x_{i}$ as the input. In summary, the concatenated vector forms the input to the unified anomaly score evaluator $\varphi$ to get an anomaly score:
$score_{i}=\varphi(\text{cat}(e_{i},z_{i},s_{i});\Theta_{\varphi})$ (11)
TABLE I: Overview of Datasets. $D$ and $N$ denote data dimension and size, respectively. $\delta$ indicates the anomaly ratio.
@@ III-E Unified Anomaly Scoring Module
III-F Training and Inference
Pre-training and Initialization.
Following [37], we pre-train an AutoEncoder on unlabeled data, initially using denoising autoencoders for layer-wise training. After this, we fine-tune the entire AutoEncoder by minimizing reconstruction loss to refine the latent representations. Once fine-tuned, the encoder is used to extract latent representations of the unlabeled data. These representations are then clustered using k-means to determine $k$ cluster centers, which are used to initialize the normal prototypes.
Training
Our model is trained in an end-to-end manner using mini-batches consisting of $b_{u}$ unlabeled samples and $b_{a}$ labeled anomalies. For each mini-batch, we compute three losses: $L_{recon}$, $L_{np}$, and $L_{score}$, as defined by Equation (3), (10), and (12), respectively. To effectively balance these three losses dynamically, we utilize a technique known as Dynamic Averaging, inspired by [38]. The total loss is calculated as follows:
$\mathcal{L_{\text{total}}}=\lambda_{1}\mathcal{L_{\text{{recon}}}}+\lambda_{2}\mathcal{L_{\text{np}}}+\lambda_{3}\mathcal{L_{\text{score}}}$ (13)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are updated per epoch based on the rate of descent of the corresponding losses.
Inference
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
This section outlines the datasets used in our experiments, covering a broad range of scenarios across domains like network security, Internet of Thing (IoT) security, financial fraud detection, and others. All 15 datasets are summarized in Table I, providing robust benchmarks for evaluating anomaly detection methods with their diverse scenarios and varying anomaly ratios. The datasets include:
  • UNSW-NB15: Network traffic records with various attack types such as DoS, Reconnaissance, Backdoor, and Analysis [39].
  • Aposemat IoT-23: Network traffic data from IoT devices, capturing both Command and Control (C&C) activities, direct attacks, DDoS, and Okiru [40].
  • Credit Card Fraud Detection: Identifies fraudulent transactions [15].
  • Vehicle Claims and Vehicle Insurance: Detects anomalies in vehicle-related insurance claims [41].
  • MAGIC Gamma Telescope: Classifies high-energy gamma particles in physics and chemistry [42].
  • Census Income: Predicts high-income individuals [43].
  • Pendigits: A pen-based handwritten digit recognition dataset [44].
  • Bank Marketing: Targets customer subscription predictions [45].
IV-A2 Baseline Methods
To construct a comprehensive comparison, we adopt several different types of baseline methods. These algorithms are grouped into two categories based on their supervision level: 1) Weakly supervised algorithms, such as RoSAS [17], PReNet [15], DevNet [16], DeepSAD [19], and FeaWAD [21], leverage limited labeled data alongside a larger set of unlabeled data, offering a direct comparison to our approach; 2) Unsupervised algorithms like DeepIForest [46], DeepSVDD [18], and iForest [47], operate without any labeled data, offering a baseline to assess the benefits of incorporating even minimal labeled data in anomaly detection. By comparing our method against these robust and diverse algorithms, we aim to demonstrate its effectiveness and flexibility in detecting anomalies with minimal labeled data.
TABLE II: Performance comparison of different methods. The upper section presents AUC-PR (primary metric) and the lower section shows AUC-ROC. The best results are highlighted in bold and the second-best results are underlined.